Three point-like charges q1 = 12 µC, q2 = -7 µC and q3 = 5 µC are located at the following positions: r1 = (0, 0, 0), r2 = (2 cm, 0, 0), and r3 = (0, 4 cm, 0), respectively. By using the Coulomb’s law (see Eq. 2.4 in the textbook) find the electric field at an arbitrary point r = (x,y,z). Then, by making use of the general result, calculate the field at (0, 0, 2 cm).
Answer #1

ที่ไ3 3 Y-1 = 0(0,๐ QV ·02-140.02 K 1118.1359K 12 =(981e E,= [55.7 re-3)「6(89.988 í-32 34.27 19 J + 24926-14 7ㄱ _a..IN4 J+ 22